Carl and Dennis are brothers. The sum of their ages is 29. In 6 years' time Carl will be the same age as Dennis was 3 years ago.


1
Expert's answer
2022-04-12T17:51:55-0400

Let Carl be "x" years old

Let Dennis be "y" years old

Then,

"\begin{cases}
x + y = 29 \\
x + 6 = y - 3
\end{cases}
\begin{cases}
x + y = 29 \\
x - y = - 9
\end{cases}
+"

"2x = 20 \\
x = 10"

Then

"y = 29 - 10 \\
y = 19"

Answer:

Carl is 10 years old

Dennis is 19 years old